OpenStudy (anonymous):

(5+y)/(5-y)=3 solve for y

OpenStudy (anonymous):

5 + y/5-y = 3 5 +y/5-y x 5-y = 3(5-y) 5 + y = 15 - 3y 5 + y -5 = 15 -3y -5 y = 10 - 3y y + 3y = 10 4y = 10 y = 2.5 sorry i don't know for sure, but i think this is what it is.

OpenStudy (campbell_st):

well first off look at the denominator and see what y can't be. it can't be 5 as this would result in a zero denominator so multiply both sides by ( 5 - y) and you get 5 + y = 3(5 - y) distribute then collect like terms and you can then solve for y hope this helps

OpenStudy (anonymous):

tnx for the help

OpenStudy (anonymous):

Yep just tested it y= 2.5 :D good luckk

OpenStudy (anonymous):

yes it works 2.5